# MR-EEGWaveNet: Multiresolutional EEGWaveNet for Seizure Detection from Long EEG Recordings

### v1

@@ Abstract
Feature engineering for generalized seizure detection models remains a significant challenge. Recently proposed models show variable performance depending on the training data and remain ineffective at accurately distinguishing artifacts from seizure data. In this study, we propose a novel end-to-end model, “Multiresolutional EEGWaveNet (MR-EEGWaveNet),” which efficiently distinguishes seizure events from background electroencephalogram (EEG) and artifacts/noise by capturing both temporal dependencies across different time frames and spatial relationships between channels. The model has three modules: convolution, feature extraction, and predictor. The convolution module extracts features through depth-wise and spatio-temporal convolution. The feature extraction module individually reduces the feature dimension extracted from EEG segments and their sub-segments. Subsequently, the extracted features are concatenated into a single vector for classification using a fully connected classifier called the predictor module. In addition, an anomaly score-based post-classification processing technique was introduced to reduce the false-positive rates of the model. Experimental results were reported and analyzed using different parameter settings and datasets (Siena (public) and Juntendo (private)). The proposed MR-EEGWaveNet significantly outperformed the conventional non-multiresolution approach, improving the F1 scores from $0.177$ to $0.336$ on Siena and $0.327$ to $0.488$ on Juntendo, with precision gains of $15.9\%$ and $20.62\%$, respectively.
@@ 1 Introduction
The performance of the MR-EEGWaveNet is evaluated with a public dataset (“Siena”) and a private dataset (“Juntendo”). While direct comparison with existing methods is difficult due to differing parameters, evaluation strategies, and datasets, we performed a comparative analysis with a recent state-of-the-art model. The results indicate that the MR-EEGWaveNet performs better in several key evaluation metrics.
@@ 2 Previous Works
Several studies have previously explored potential solutions to the seizure detection problem [18, 19, 20, 16]. They proposed several models and extracted features to distinguish seizure and nonseizure events from long recordings of EEG. In EEG research, the calculated features used to separate brain activity are commonly categorized into three groups: complexity, continuity, and connectivity [21]. In epilepsy research, complexity features (such as entropy and fractal-based methods) measure the irregularity in the data to detect seizure activity in EEG. Recently, different types of entropy have been proposed in the seizure detection task [22, 23]. Continuity-based features, such as band power ($\delta$, $\theta$, $\alpha$, $\beta$, and $\gamma$), median frequency, diffuse slowing, and burst suppression [24], calculate the stability and persistence of EEG signals over time. Connectivity features, such as coherence, phase-locking value (PLV), and mutual information, assess the relationship and interaction between EEG channels that cover different regions of the brain [21]. Furthermore, a combination of time-, frequency-, and entropy-based features is applied with a random forest classifier to detect seizures in EEG [25].
@@
However, the nonlinear and nonstationary nature of EEG presents a significant challenge in extracting features that consistently perform well over time. In general, EEG signals exhibit substantial inter-subject variability, and the features extracted from them are often highly subject-specific. In recent years, researchers have been actively developing data-driven approaches, such as end-to-end deep learning models to overcome feature engineering challenges in EEG [20, 16]. Convolutional neural networks (CNN) and bidirectional long short-term memory are well-known models that are used together to detect seizures, as proposed in [20]. Using a discrete wavelet transform (DWT), the EEG signals are decomposed into subbands, and the features are extracted using CNNs. Subsequently, the false detection rate (FDR) and sensitivity were improved by smoothing and collar techniques. By transforming EEG into plot images, a CNN-based model for seizure detection was designed in [31]. The model was improved with the addition of a patient-specific autoencoder (AE), which introduced a new label, “nonseizure-but-abnormal,” to the existing binary labels (“seizure” and “nonseizure”) [32]. Recently, graph-based analysis has also gained popularity in seizure detection. In [33], the short-time Fourier transform (STFT) of EEG signals is transformed into graph data and then classified with Google-Net CNN models. In [34], the EEG signal was first transformed into a spectrogram using the Stockwell transform. A Linear Graph Convolution Network (GCN) was used for feature selection, followed by DenseNet-based classification. “EEGWaveNet” is a deep learning model to process EEG signals introduced in [16]. It can capture both time-related and channel-related features that are suitable for classifying EEG signals compared to other models. However, implementing deep learning-based models in the medical domain (specifically seizure detection) has faced several challenges. Due to the complex and subject-dependent nature of seizures, these techniques have faced challenges, such as extreme data imbalance, complex interpretation, and the need for large amounts of training data to fine-tune hyperparameters.
@@ 3.1 Public Dataset
A publicly available EEG dataset was employed. The dataset was recorded at the Unit of Neurology and Neurophysiology of the University of Siena [39]; hereafter, it will be referred to as “Siena”. The dataset comprised recordings of 14 patients, including nine males (ages 25–71) and five females (ages 20–58). The International 10-20 standard was used for EEG recordings with the $512$ Hz sampling frequency. The EEG recordings were downsampled to $500$ Hz for the experiment. The dataset was diagnosed and labeled (seizure events) by an expert clinician according to the International League Against Epilepsy criteria [40]. This study was conducted using EEG recordings from 19 commonly used channels: Fp1, Fp2, F7, F8, F3, F4, T3, T4, T5, T6, C3, C4, P3, P4, Fz, Cz, Pz, O1, and O2. Detailed information and statistics of the dataset are summarized in Table 1. The dataset is publicly available and can be accessed in [40].
@@ 3.2 Private Dataset
In this study, a private EEG dataset recorded at Juntendo University Hospital in Japan was employed. The ethics committees of Juntendo University Hospital and Tokyo University of Agriculture and Technology approved the research. The dataset consists of scalp EEG recordings from multiple epileptic patients, with several hours of data collected for each. Specifically, 21 patients’ data (fourteen males and seven females), ranging in age from 1 to 42 years, were used in the experiment. The electrode placement followed the International 10-20 system, with a sampling frequency of 500 Hz. The number of selected channels was 19, the same as the Siena EEG dataset. The dataset contains different types of seizures, such as generalized, motor, nonmotor, focal, non-focal, tonic-clonic, behavioral arrest, among others. The brief data statistics are shown in Table 2. The $10$-s long nonseizure and seizure EEG segment from the dataset is shown in Figure 1.
@@ A Convolution Module
Input ($C$, $N$) Linear
Conv1D kernel 2, stride 2, group $C$ ($C$, $N$/2) Linear
Conv1D kernel 2, stride 2, group $C$ ($C$, $N$/4) Linear
Conv1D kernel 2, stride 2, group $C$ ($C$, $N$/8) Linear
Conv1D kernel 2, stride 2, group $C$ ($C$, $N$/16) Linear
Conv1D kernel 2, stride 2, group $C$ ($C$, $N$/32) Linear
Conv1D kernel 2, stride 2, group $C$ ($C$, $N$/64) Linear
@@
Activation Leaky ReLU
Conv1D (32$\times$4) (32, $N$/$62^{k}-$ 6)
@@ B Feature Extraction Module
Fully Connected 64 Leaky ReLU
Fully Connected 32 Sigmoid
@@ C Predictor Module
Fully Connected 64 Leaky ReLU
@@
Figure 2: Classification process of MR-EEGWaveNet model a) Convolution Module, and b) Overall MR-EEGWaveNet Architecture (Here, $10$ s, and $2$ s represent $10$ seconds, and $2$ seconds long multichannel EEG signal)
@@
The MR-EEGWaveNet model contains three main modules: convolution, feature extraction, and predictor. The “convolution” module has a series of convolutional layers and pooling operations to process EEG signals, and is divided into two parts: a multi-scale convolution module and a spatio-temporal convolution module. The “multi-scale convolution” module has six temporal convolutional layers, each reducing the signal sequence length while maintaining channel information. The output of each temporal convolutional layer feeds to the next convolutional layer (except the last convolutional layer). The “spatio-temporal convolution” module performs pooling operations across multiple blocks to extract features from all channels. In each block, there are three operational components: convolutional layers, batch normalization, and LeakyReLU activation. For the EEGWaveNet architecture, we used two blocks sequentially to calculate features. Here, the inputs are the outputs of the temporal convolutional layers from “multi-scale convolution” (except the first temporal convolutional layer). Thus, five “spatio-temporal convolution” modules were used in the proposed model, each producing an output of $1\times 32$. Subsequently, they were concatenated to form a vector of dimension $1\times 160$.
@@
Thereafter, the “feature extraction” module shrank the dimension of the feature from $160$ to $32$ with two fully connected layers with LeakyReLU activation. Thus, the final output of modules A and B for an EEG segment was $32$. The same process applies to the sub-segments. For example, if the proposed model has multiresolutional parameters $10$ s and $2$ s, then it will calculate one feature vector ($1\times 32$) for $10$ s and five feature vectors corresponding to five sub-segments of $2$ s (each has the same dimension of $1\times 32$). After concatenation, the final size of the feature vector will be $1\times 192$. Here, the list of multiresolutional parameters defines the architecture of the model. Figure 2 shows the architecture for the parameter list “$10$ s, $2$ s”. In this case, the model subdivided the $10$-s EEG signal into $2$ s segments; in total, six EEG signals (one $10$ s and five $2$ s multichannel EEG signals) were generated. Each signal was fed to modules A through B to calculate the features of each of them. Finally, the concatenated feature vector was fed into the predictor module for classification. The predictor module contains a fully connected classifier that maps the features to the output classes. The class probabilities were computed using the log-softmax function. The convolution module and the detailed architecture of the MR-EEGWaveNet for the above example are shown in Figures 2(a) and 2(b), respectively.
The main difference between EEGWaveNet and MR-EEGWaveNet architecture is that MR-EEGWaveNet extracts features both from the EEG segments and their sub-segments using modules A and B (the convolution and feature extraction module). Thus, the EEG segment was segmented into multiple levels based on the parameters. The architecture of MR-EEGWaveNet is shown in Table 3.
@@ 4.2 Post-Classification Processing
To further improve the performance of the proposed architecture, we applied an anomaly score-based post-classification processing technique to reduce the false detection rate of seizures. Therefore, we calculated the anomaly scores of the EEG segments, $a_{i}$ ($i$ is the segment index), with one of the state-of-the-art anomaly detection methods named ECOD: Unsupervised Outlier Detection Using Empirical Cumulative Distribution Functions [41]. Finally, the following rule was applied for deciding a segment as seizure or nonseizure based on the anomaly score of the EEG segments:
@@
$L_{i}$ =$=\begin{cases}1,&\text{seizure , if }a_{i}>\mu_{a}\\ 0,&\text{nonseizure}.\end{cases}$ (2)
@@ 5 Experiment
The raw EEG data from each patient were preprocessed and segmented into equal length, $W$ (window length in seconds), using a sliding window approach with no overlap. Each segment of EEG has a dimension of $C\times N$, where $C$ is the number of channels, and $N$ is the number of time points of a single channel. Let $F_{s}$ be the sampling frequency, then $N=WF_{s}$.
For EEGWaveNet, the feature dimension is $F$=$32$, whereas in the case of MR-EEGWaveNet, it depends on the multiresolution parameter list $D$ (such as [$10$ s, $5$ s, $2$ s]). The number of parameters and the computational complexity of the models depend on $D$, $F$, and $W$. Here, the concatenated feature length $K$ depends on them and can be calculated as:
@@
[…]
@@
True Positive Rate $\text{True Positive Rate or}\;\mathrm{Recall}$ $=\frac{TP}{TP+FN}$ (5)
@@
$\mathrm{FPR}$ $=\frac{FP}{FP+TN}$ (8)
where $TP$, $TN$, $FP$, and $FN$ represent the total number of correctly detected positive, correctly detected negative, incorrectly detected positive, and incorrectly detected negative EEG segments, respectively. Moreover, the Receiver Operating Characteristic–Area Under the Curve (ROC-AUC) is a standard metric used to evaluate the ability of a binary classification model to differentiate between two classes. The ROC curve plots the True Positive Rate (TPR) (also known as recall) against the FPR at various classification thresholds. The Area Under the Curve (AUC) measures the entire area beneath the ROC curve and quantifies the overall performance of a model. The AUC is a single value that ranges from 0 to 1. An AUC of 1.0 indicates perfect classification, while an AUC of 0.5 does not reflect discriminative ability, the same as random guessing. Generally, a classifier with a higher AUC is considered to be more effective in binary classification tasks. Finally, the Det. Ratio represents the seizure detection ratio from the outcome of a model.
The models were evaluated with a LOSO validation approach, where one patient was designated for testing and the rest for training. In the case of preparing a dataset for training, we adopted the following two techniques to overcome the data imbalance problem:
  1. I.
    Seizure data: Oversampling; the seizure segments (within the expert’s notation of the seizure label) are 80%-overlapped between consecutive segments.
  2. II.
    Nonseizure data: Undersampling; the nonseizure segments are randomly selected from interictal EEG [42], the ratio between seizure and nonseizure is 1:2. This selection ratio is maintained for all recording sessions individually (that is, for each recording file).
For the test set, the recordings were segmented with a non-overlapping sliding window approach with a fixed value of $W$, where each segment was assigned to a label either 0 (nonseizure, negative class) or 1 (seizure, positive class) based on the ground truth label provided by an expert.
We used the following parameter values to train each model: batch size: $32$, number of training epochs: $300$, and learning rate: $10^{-3}$. The cross-entropy loss function was employed with balanced weights of $0.75$ and $1.5$ for the nonseizure and seizure classes, respectively. A warm start strategy with an early stop was used, where training continues until the validation loss stops improving for consecutive epochs (patience=$30$). For each iteration, the training and validation sets were randomly sampled, and the test set remained fixed across all runs. Each experiment was conducted three times to get a stable result. For each patient, performance metrics were averaged across the three runs, that is, the final score for a patient was computed as the mean of their results over the three iterations.

### v2

@@ Abstract
Feature engineering for generalized seizure detection models remains a significant challenge. Recently proposed models show variable performance depending on the training data and remain ineffective at accurately distinguishing artifacts from seizure data. In this study, we propose a novel end-to-end model, “Multiresolutional EEGWaveNet (MR-EEGWaveNet),” which efficiently distinguishes seizure events from background electroencephalogram (EEG) and artifacts/noise by capturing both temporal dependencies across different time frames and spatial relationships between channels. The model has three modules: convolution, feature extraction, and predictor. The convolution module extracts features through depth-wise and spatio-temporal convolution. The feature extraction module individually reduces the feature dimension extracted from EEG segments and their sub-segments. Subsequently, the extracted features are concatenated into a single vector for classification using a fully connected classifier called the predictor module. In addition, an anomaly score-based post-classification processing technique is introduced to reduce the false-positive rates of the model. Experimental results are reported and analyzed using different parameter settings and datasets (Siena (public) and Juntendo (private)). The proposed MR-EEGWaveNet significantly outperformed the conventional non-multiresolution approach, improving the F1 scores from $0.177$ to $0.336$ on Siena and $0.327$ to $0.488$ on Juntendo, with precision gains of $15.9\%$ and $20.62\%$, respectively.
@@ 1 Introduction
The performance of the MR-EEGWaveNet is evaluated with a public dataset (“Siena”) and a private dataset (“Juntendo”). While direct comparison with existing methods is difficult due to differing parameters, evaluation strategies, and datasets, we performed a comparative analysis with recent state-of-the-art models. The results indicate that the MR-EEGWaveNet performs better in several key evaluation metrics.
@@ 2 Previous Works
Several studies have previously explored potential solutions to the seizure detection problem [18, 19, 20, 16]. They proposed several models and extracted features to distinguish seizure and nonseizure events from long recordings of EEG. In EEG research, the calculated features used to separate brain activity are commonly categorized into three groups: complexity, continuity, and connectivity [21]. In epilepsy research, complexity features (such as entropy and fractal-based methods) measure the irregularity in the data to detect seizure activity in EEG. Recently, different types of entropy have been proposed for the seizure detection task [22, 23]. Continuity-based features, such as band power ($\delta$, $\theta$, $\alpha$, $\beta$, and $\gamma$), median frequency, diffuse slowing, and burst suppression [24], calculate the stability and persistence of EEG signals over time. Connectivity features, such as coherence, phase-locking value, and mutual information, assess the relationship and interaction between EEG channels that cover different regions of the brain [21]. Furthermore, a combination of time-, frequency-, and entropy-based features has been applied with a random forest classifier to detect seizures in EEG [25].
@@
However, the nonlinear and nonstationary nature of EEG presents a significant challenge in extracting features that consistently perform well over time. In general, EEG signals exhibit substantial inter-subject variability, and the features extracted from them are often highly subject-specific. In recent years, researchers have been actively developing data-driven approaches, such as end-to-end deep learning models to overcome feature engineering challenges in EEG [20, 16]. Convolutional neural networks (CNN) and bidirectional long short-term memory are well-known models that are used together to detect seizures, as proposed in [20]. Using a discrete wavelet transform, the EEG signals were decomposed into subbands, and the features were extracted using CNNs. Subsequently, the false detection rate and sensitivity were improved by smoothing and collar techniques. By transforming EEG into plot images, a CNN-based model for seizure detection was designed in [31]. The model was improved with the addition of a patient-specific autoencoder, which introduced a new label, “nonseizure-but-abnormal,” to the existing binary labels (“seizure” and “nonseizure”) [32]. Recently, graph-based analysis has also gained popularity in seizure detection. In [33], the short-time Fourier transform (STFT) of EEG signals was transformed into graph data and then classified with Google-Net CNN models. In [34], the EEG signal was first transformed into a spectrogram using the Stockwell transform. A Linear Graph Convolution Network (GCN) was used for feature selection, followed by DenseNet-based classification. “EEGWaveNet” is a deep learning model to process EEG signals introduced in [16]. It can capture both time-related and channel-related features that are suitable for classifying EEG signals compared to other models. However, implementing deep learning-based models in the medical domain (specifically seizure detection) has faced several challenges. Due to the complex and subject-dependent nature of seizures, these techniques have faced challenges, such as extreme data imbalance, complex interpretation, and the need for large amounts of training data to fine-tune hyperparameters.
EEGNet, proposed in [35], is a widely adopted model specifically developed for brain-computer interface tasks. Through the use of temporal, depthwise, and separable convolutions, the model captures key frequency elements, spatial features specific to those frequencies, and temporal patterns. In contrast, “EEG Conformer” adopts a compact convolution-attention architecture [36]. It captures short-term temporal and spatial features using convolutional layers, while self-attention models long-term global temporal dependencies. This design enables a unified and robust framework for EEG classification.
The success of large language models in other domains has also inspired researchers to develop EEG-based foundation models recently [37, 38, 39]. Biosignal Transformer (BIOT) [37] preprocesses biosignals into tokenized sequences and employs a linear transformer to capture temporal interactions. Its encoder supports supervised learning, pre-training, and fine-tuning on both complete and incomplete data across various biosignal tasks. In [38], the authors proposed a Large Brain Model (LaBraM) to overcome challenges such as varying electrode counts, low signal-to-noise ratios, and more. The LaBraM model segments EEG signals into fixed-length patches, applies temporal encoding, adds temporal and spatial embeddings, and then processes the sequence with a Transformer encoder using patch-wise attention for final feature extraction. Recently, a brain foundation model named Criss-Cross Brain Foundation Model (CBraMod) has been proposed [39]. It employs a criss-cross transformer backbone to independently model spatial and temporal dependencies in EEG signals via parallel attention. Additionally, It uses an asymmetric conditional positional encoding scheme that encodes EEG patch positions and easily adapts to diverse EEG formats.
@@ 3.1 Public Dataset
A publicly available EEG dataset was employed. The dataset was recorded at the Unit of Neurology and Neurophysiology of the University of Siena [44]; hereafter, it will be referred to as “Siena”. The dataset comprised recordings of 14 patients, including eight males (ages 25–71) and six females (ages 20–58). The International 10-20 standard was used for EEG recordings with the $512$ Hz sampling frequency. The EEG recordings were downsampled to 500 Hz for the experiment. The dataset was diagnosed and labeled (seizure events) by an expert clinician according to the International League Against Epilepsy criteria [45]. This study was conducted using EEG recordings from 19 commonly used channels: Fp1, Fp2, F7, F8, F3, F4, T3, T4, T5, T6, C3, C4, P3, P4, Fz, Cz, Pz, O1, and O2. Detailed information and statistics of the dataset are summarized in Table 1. The dataset is publicly available and can be accessed in [45].
@@ 3.2 Private Dataset
In this study, a private EEG dataset recorded at Juntendo University Hospital in Japan was employed. The ethics committees of Juntendo University Hospital and Tokyo University of Agriculture and Technology approved the research. The dataset consists of scalp EEG recordings from multiple epileptic patients, with several hours of data collected for each. Specifically, 21 patients’ data (thirteen males and eight females), ranging in age from 1 to 42 years, were used in the experiment. The electrode placement followed the International 10-20 system, with a sampling frequency of 500 Hz. The number of selected channels was 19, the same as the Siena EEG dataset. The dataset contains different types of seizures, such as generalized, motor, nonmotor, focal, non-focal, tonic-clonic, behavioral arrest, among others. The brief data statistics are shown in Table 2. The $10$-s long nonseizure and seizure EEG segment from the dataset is shown in Figure 1.
@@ A Convolution Module
Input ($C$, NN) Linear
Conv1D kernel 2, stride 2, group $C$ ($C$, NN/2) Linear
Conv1D kernel 2, stride 2, group $C$ ($C$, NN/4) Linear
Conv1D kernel 2, stride 2, group $C$ ($C$, NN/8) Linear
Conv1D kernel 2, stride 2, group $C$ ($C$, NN/16) Linear
Conv1D kernel 2, stride 2, group $C$ ($C$, NN/32) Linear
Conv1D kernel 2, stride 2, group $C$ ($C$, NN/64) Linear
@@
Activation LeakyReLU(0.01)
[…]
@@ B Feature Extraction Module
Fully Connected 64 LeakyReLU(0.01)
Fully Connected 32
Normalization 32
@@ C Predictor Module
Fully Connected 64 LeakyReLU(0.01)
@@
Figure 2: Classification process of MR-EEGWaveNet model (a) Convolution module, and (b) Overall MR-EEGWaveNet architecture (Here, $10$ s, and $2$ s represent $10$-second, and 22-second long multichannel EEG signal)
@@
The architecture of MR-EEGWaveNet is shown in Table 3. The MR-EEGWaveNet model contains three main modules: convolution, feature extraction, and predictor. The “convolution” module has a series of convolutional layers and pooling operations to process EEG signals, and is divided into two parts: a multi-scale convolution module and a spatio-temporal convolution module. The “multi-scale convolution” module has six temporal convolutional layers, each reducing the signal sequence length while maintaining channel information. The output of each temporal convolutional layer feeds to the next convolutional layer (except the last convolutional layer). The “spatio-temporal convolution” module performs pooling operations across multiple blocks to extract features from all channels. In each block, there are three operational components: convolutional layers, batch normalization, and LeakyReLU activation. For the MR-EEGWaveNet architecture, we used two blocks sequentially to calculate features. Here, the inputs are the outputs of the temporal convolutional layers from “multi-scale convolution” (except the first temporal convolutional layer). Thus, five “spatio-temporal convolution” modules are used in the proposed model, each producing an output of $1\times 32$. Subsequently, they are concatenated to form a vector of dimension $1\times 160$.
@@
Thereafter, the “feature extraction” module shrank the dimension of the feature from $160$ to $32$ with two fully connected layers, followed by normalization. Thus, the final output of modules A and B for an EEG segment is $32$. The same process applies to the sub-segments. For example, if the proposed model has multiresolutional parameters $10$ s and $2$ s, then it will calculate one feature vector ($1\times 32$) for $10$ s and five feature vectors corresponding to five sub-segments of $2$ s (each has the same dimension of $1\times 32$). After concatenation, the final size of the feature vector will be $1\times 192$. Here, the list of multiresolutional parameters defines the architecture of the model. Figure 2 shows the architecture for the parameter list “$10$ s, $2$ s”. In this case, the model subdivides the $10$-s EEG signal into $2$ s segments; in total, six EEG signals (one $10$ s and five $2$ s multichannel EEG signals) are generated. Each signal is fed to modules A through B to calculate the features of each of them. Finally, the concatenated feature vector is fed into the predictor module for classification. The predictor module contains a fully connected classifier that maps the features to the output classes. The class probabilities are computed using the log-softmax function. The convolution module and the detailed architecture of the MR-EEGWaveNet for the above example are shown in Figures 2(a) and 2(b), respectively. The main difference between EEGWaveNet and MR-EEGWaveNet architecture is that MR-EEGWaveNet extracts features both from the EEG segments and their sub-segments using modules A and B (the convolution and feature extraction module). Thus, the EEG segment is segmented into multiple levels based on the parameters.
@@ 4.2 Post-Classification Processing
To further improve the performance of the proposed architecture, we applied an anomaly score-based post-classification processing technique to reduce the false detection rate of seizures. Therefore, we calculated the anomaly scores of the EEG segments, $a_{i}$ (ii is the segment index), using one of the state-of-the-art anomaly detection methods, ECOD: Unsupervised Outlier Detection Using Empirical Cumulative Distribution Functions [46]. Finally, the following rule is applied to decide whether a segment is classified as a seizure or nonseizure based on the anomaly score of the EEG segments:
@@
$L_{i}$ $=\begin{cases}1,&\text{seizure, if }a_{i}>\mu_{a}\\ 0,&\text{nonseizure}.\end{cases}$ (2)
@@ 5 Experiment
The raw EEG data from each patient were preprocessed and segmented into equal length, $W$ (window length in seconds), using a sliding window approach with no overlap. Each segment of EEG has a dimension of $C\times N$, where $C$ is the number of channels, and $N$ is the number of time points of a single channel. In the MR-EEGWaveNet (see Table 3), $C=19$, and $k$ is the order of layers in Module A(II). Let $F_{s}$ be the sampling frequency, then $N=WF_{s}$.
For EEGWaveNet, the feature dimension is FF=$32$, whereas in the case of MR-EEGWaveNet, it depends on the multiresolution parameter list $D$ (such as [$10$ s, $5$ s, $2$ s]). The number of parameters and the computational complexity of the models depend on $D$, $F$, and $W$. Here, the concatenated feature length $K$ depends on them and can be calculated as:
@@
where $\left\lfloor\cdot\right\rfloor$ denotes the floor function producing the largest integer less than or equal to the argument.
@@
Moreover, to evaluate the contribution of individual components within the MR-EEGWaveNet architecture, we conducted an ablation study using two simplified variants: MR-EEGWaveNet-Abl1 and MR-EEGWaveNet-Abl2. We selected the MR-EEGWaveNet-2 model as the baseline, which guided both the development of ablation variants and the experimental comparison. The two ablation models are defined as follows:
  1. i)
    MR-EEGWaveNet-Abl1 excludes the feature extraction module from the baseline model.
  2. ii)
    MR-EEGWaveNet-Abl2 removes the $10$-second input pipeline, retaining only the $2$-second sub-segments stream.
The detailed architectures of the ablated models, MR-EEGWaveNet-Abl1 and MR-EEGWaveNet-Abl2, are provided in Appendix A, in Figures A1 and A2, respectively.
In the case of calculating anomaly scores using ECOD, we applied it directly to the preprocessed EEG segments rather than to learned embeddings. For each segment, all channels were concatenated into a single feature vector by flattening the multichannel signal ($C$ channels × $N$ time samples per channel $\rightarrow$ $C\times N$-dimensional vector). The flattened signals were z-score normalized (standardized to have zero mean and unit variance) before further processing. ECOD was then applied to these fixed-length vectors to compute anomaly scores for each segment in a recording, which were classified as anomalous based on Equations 1 and 2. The ECOD implementation can be found at [47, 48].
@@
$\mathrm{True\ Positive\ Rate\ (Recall)}$ $=\frac{TP}{TP+FN}$ (5)
@@
$\mathrm{F1\ Score}$ $=2\times\frac{\mathrm{Precision}\times\mathrm{Recall}}{\mathrm{Precision}+\mathrm{Recall}}$ (7)
$\mathrm{False\ Positive\ Rate\ (FPR)}$ $=\frac{FP}{FP+TN}$ (8)
Where $TP$, TNTN, FPFP, and $FN$ denote the total number of true positives, true negatives, false positives, and false negatives, respectively, based on the detection of EEG segments. Moreover, the Receiver Operating Characteristic–Area Under the Curve (ROC-AUC) is a standard metric used to evaluate the ability of a binary classification model to differentiate between two classes. The ROC curve plots the True Positive Rate (TPR) (also known as recall) against the FPR at various classification thresholds. The Area Under the Curve (AUC) measures the entire area beneath the ROC curve and quantifies the overall performance of a model. The AUC is a single value that ranges from 0 to 1. An AUC of 1.0 indicates perfect classification, while an AUC of 0.5 does not reflect discriminative ability, the same as random guessing. Generally, a classifier with a higher AUC is considered to be more effective in binary classification tasks. Finally, the Det. Ratio represents the seizure detection ratio from the outcome of a model.
The models were evaluated with a LOSO validation approach, where in each fold, one patient was designated for testing and the remaining patients were used for training, as illustrated in Figure 3. In the case of preparing a dataset for training, we adopted the following two techniques to overcome the data imbalance problem:
  1. I.
    Seizure data: Oversampling; the seizure segments (within the expert’s notation of the seizure label) are $80\%$-overlapped between consecutive segments.
  2. II.
    Nonseizure data: Undersampling; the nonseizure segments (non-overlapping) are randomly selected from interictal EEG [49], the ratio between seizure and nonseizure is 1:2.
For the test set, the recordings were segmented using a non-overlapping sliding window approach with a fixed window size $W$, where each segment was assigned a label either 0 (nonseizure, negative class) or $1$ (seizure, positive class) based on the ground truth label provided by an expert (similar to [25]).
Figure 3: Illustration of Leave-One-Subject-Out (LOSO) cross-validation scheme
We used the following parameter values to train each model: batch size: $32$, number of training epochs: $300$, and Adam optimizer with learning rate: $10^{-3}$. The cross-entropy loss function was employed with balanced weights of $0.75$ and $1.5$ for the nonseizure and seizure classes, respectively. The optimizer was reinitialized at the beginning of each training epoch, effectively resetting the optimization state while retaining the current model parameters. An early stopping criterion was employed, where training was stopped if the validation loss did not improve for $30$ consecutive epochs (patience = $30$). For each training epoch, nonseizure samples were randomly selected at twice the number of seizure samples (i.e., a 1:2 seizure-to-nonseizure ratio). This balanced subset is then randomly split: $70\%$ for training and $30\%$ for validation. Each experiment was repeated three times to ensure stability of the results. For each patient, the performance metrics were averaged across the three runs; that is, the final score for a patient was computed as the mean of their results over the three iterations.
@@
[…]